# Structural Change, Employment, and Inequality in Europe: an Economic Complexity Approach

### v1

@@ 2.1 Structural change, Fitness and Complexity
In order to measure the Fitness of countries, and more importantly for our purposes, the Complexity of European industries, we rely on the Fitness and Complexity algorithm proposed by Tacchella et al. (2012):
$\widetilde{F}_{c}^{(n)}=\sum_{i}M_{ci}Q_{i}^{(n-1)}&&F_{c}^{(n)}=\frac{\widetilde{F}_{c}^{(n)}}{\langle\widetilde{F}_{c}^{(n)}\rangle_{g}}\\ \\ \widetilde{Q}_{i}^{(n)}=\frac{1}{\sum_{c}M_{ci}\frac{1}{F_{c}^{(n)}}}&&Q_{i}^{(n)}=\frac{\widetilde{Q}_{i}^{(n)}}{\langle\widetilde{Q}_{i}^{(n)}\rangle_{a}}\end{array}\right.$ (1)
with $F^{(0)}_{c}=1$ and $Q^{(0)}_{i}=1$.
@@
The algorithm is fed with a country-industry employment matrix constructed by cross-tabulating industrial employment data across European countries, adequately binarised [^1] The binary specialisation matrix – which identifies industries in which countries are specialised – is obtained using an entropy maximising null model (the Bipartite Weighted Configuration Model) and implementing a method to ensure the comparability of fitness and complexity values over time. A more in-depth explanation of the construction of the matrix are explained in section 4.1, from which we obtain the values Fitness (first row of equation 1) and Complexity (second row of equation 1). In practical terms, the Fitness $F$ of country $c$ at each time period corresponds to the sum of the Complexities $Q$ of the industries $i$ in which country $c$ holds a comparative advantage, the latter being defined by the row sums of the binary specialisation matrix $M_{c,i}$. In turn, the complexity $Q$ of industry $i$ corresponds to the inverse of its ubiquity (the column sums of $M_{c,i}$) weighted by the inverse of the Fitness of the countries which have that industry in their productive portfolio. In short, more complex industries are the least ubiquitous ones – proxying for the rareness of the capabilities required by those – considering also how "fit" are the countries that hold a specialisation in that industry. The algorithm is iterated $n$ times, until it converges. At each step, Fitness and Complexity values are normalised by dividing them by the respective average values.
A feature that makes Fitness our preferred measure of knowledge intensity of the productive structure is that it considers the cumulative nature of capabilities, giving importance to the diversification of countries (weighted by the complexity of their industries). This feature is absent in competing methods, such as the Economic Complexity Index (ECI) (Hausmann & Hidalgo 2009), which has been shown to measure country fitness as the average complexity of the economic activities in which a country holds a revealed comparative advantage (Cristelli et al. 2013).
@@ 2.2 Labour-Weighted Fitness and decomposition
The starting point to construct a measure of structural change that captures the reallocation of labour across industries is to manipulate the Fitness equation (the first row of equation 1) in such a way as to weight the sum of the complexities $Q_{i}$ by their labour share. Thus, the Labour-Weighted Fitness can be expressed as:
@@
Where $\theta_{i,t}$ is the labour share of industry $i$ at time $t$ for any given country. The resulting LWF is therefore an index of the Fitness of countries, where more weight is given to industries that are not only complex, but which also absorb a larger employment share, indicating the presence of knowledge-intensive capabilities across a larger portion of the workforce.
To extract our measure of structural change, we now propose a structural decomposition of LWF inspired by the productivity decompositions adopted in the structuralist literature (McMillan et al. 2014, 2017, de Vries et al. 2015), which originates from the original work of Fabricant (1942). We decompose $\Delta LWF_{t}$ – the variation of LWF between time $t$ and $t-k$, with $k\leq T$ – in within and between sector components, as follows:
@@
$(=)\ \underbrace{\sum_{i=1}^{n}\theta_{i,t-k}\Delta Q_{it}}_{\text{WITHIN TERM}}+\underbrace{\sum_{i=1}^{n}Q_{it}\Delta\theta_{i,t}}_{\text{BETWEEN TERM}}$ (6)
After manipulation, Equation 3 yields two distinct terms, described in Equation 6:
  • $\sum_{i=1}^{n}\theta_{i,t-k}\Delta Q_{it}$ is the within-sector term of structural change (the contribution of increasing complexity at the sectoral level to the change in labour-weighted Fitness)
@@
The between component, given by $\sum_{i=1}^{n}Q_{it}\Delta\theta_{i,t}$, identifies the measure of structural change of interest. This will be used to explain the observed trends in wage inequality, employment growth, and functional distribution of income.
@@ 3 Data
3.1 EUROSTAT’s Structural Business Survey
The main data source used in this paper is the Eurostat Structural Business Statistics (SBS). The SBS describe the structure and performance of industrial sectors in the in European countries. They cover the ‘business economy’ (NACE Rev. 2 sections B to N and division 95) which includes: i) Industry; ii) Construction; iii) Distributive trades; and iv) Services. The data on industries is aggregated at the 4 digit level of the NACE classification, for a total of 218 NACE industries.
SBS are based upon data for enterprises, which is obtained for national statistical offices and aggregated at the industry level based on the enterprise’s principal activity. Businesses in financial services (NACE Rev. 2 Section K) are excluded because of their specific characteristics and the unavailability of information in this industry, often due to confidentiality issues. In addition, SBS does not cover agriculture, forestry and fishing (NACE A), public administration (NACE O) and non-market services such as education (NACE P) and health (NACE Q).
As the analysis conducted here relies on patterns of industrial employment, we use the SBS data to describe the employment share of 4 digit NACE industries. To do so, we extract the SBS indicator that quantifies the number of persons employed [^2] Following the Eurostat’s SBS definition, this indicator ”is defined as the total number of persons who work in the observation unit (inclusive of working proprietors, partners working regularly in the unit and unpaid family workers), as well as persons who work outside the unit who belong to it and are paid by it (e.g. sales representatives, delivery personnel, repair and maintenance teams). It excludes manpower supplied to the unit by other enterprises, persons carrying out repair and maintenance work in the enquiry unit on behalf of other enterprises, as well as those on compulsory military service.” Source: https://ec.europa.eu/eurostat/cache/metadata/en/sbs_esms.htm#meta_update1663843213928 in each industry. In order to maximise data availability, we restrict the analysis to the 27 countries that were part of the European Union until 2018, observed between 2010 and 2018.
@@ 3.1 EUROSTAT’s Structural Business Survey
3.2 ARDECO
The Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy (ARDECO) [^3] The ARDECO dataset is available, free of charge, from the European Commission’s website: https://knowledge4policy.ec.europa.eu/territorial/ardeco-database_en, is a dataset curated and maintained by the Commission’s Joint Research Centre. It includes information on European countries and regions at different levels of disaggregation. While the dataset is based on the Eurostat’s SBS illustrated in the previous section, the data is complemented – where needed – with national and international sources. The data covers different topics, such as population, employment, labour costs, domestic product and capital formation. For the purpose of the analysis carried out in this paper, indicators on population, GDP per capita (constant prices), employment growth, compensation of employees (at constant prices) and gross value added (at constant prices) have been selected. The last two indicators are used to construct a measure of labour share of the economy, computed as the ratio of employee’s aggregate compensation over gross value added, in each country and year.
3.3 ILO Statistics on Labour Income and Inequality
The data on wage inequality is obtained from the International Labour Organisation’s Statistics on Labour Income and Inequality. We focus on data regarding the labour income distribution [^4] Available at: https://www.ilo.org/shinyapps/bulkexplorer46/?id=LAP_2LID_QTL_RT_A, which despite being based on modelled estimates, it provides information on average wages at all deciles of the labour income distribution across the countries in the sample, for the entire time period covered (2010 – 2018). In particular, we use information on labour income at the ninth, fifth and first deciles to produce wage ratios between ninth and first, fifth and first, and ninth and fifth deciles to capture the degree of inequality in labour income at different points of the distribution.
3.4 World Development Indicators
@@ 4 Methods
4.1 Specialisation matrix using BiWCM and dummy fitness
Economic Complexity applications frequently involve the need to determine the areas in which individuals or entities possess a comparative advantage. This categorization is established using a tagging scheme, which is computed using the Revealed Comparative Advantage (RCA) index, originally formulated by Balassa (Balassa 1965). In the context of our current study, the entities being assessed are countries, and the areas of specialization are represented by the number of workers in various industries. The RCA index quantifies this specialization by calculating the ratio of a country’s share of workers in a particular industry to the global share of that same sector. To express this mathematically, the RCA is:
$[RCA]_{c,i}=\frac{L_{c,i}}{\sum_{i^{\prime}}L_{c,i^{\prime}}}\left/\frac{\sum_{c^{\prime}}L_{c^{\prime},i}}{\sum_{c^{\prime},i^{\prime}}L_{c^{\prime},i^{\prime}}}\right.=\frac{L_{c,i}}{E[L]_{c,i}},\quad E[L]_{c,i}=\frac{(\sum_{i^{\prime}}L_{c,i^{\prime}})(\sum_{c^{\prime}}L_{c^{\prime},i})}{\sum_{c^{\prime},i^{\prime}}L_{c^{\prime},i^{\prime}}}$ (7)
$M_{c,i}$ is the matrix used for example in eq. (1) and is the binary filter of the RCA associating 1 when the RCA value is above the unity, otherwise it is zero.
The primary rationale for utilizing the RCA index is to establish a measure that isn’t contingent on the relative strength of diverse countries and industries, which are inherently incomparable in terms of size. Balassa’s RCA, specifically, contrasts the actual labour value of a country with an anticipated value derived under the assumption of equalising country and sector sizes (Vollrath 1991, Krantz et al. 2018). In this interpretation, RCA is a ratio between the actual number of workers and the expected output of an average country of the same size, given the knowledge of only the total labour strength, as made explicit in the second part of eq. (7). However, it’s essential to acknowledge that RCA carries inherent bias due to the heterogeneity of the country and industrial strength, resulting in correlations in the estimation of average country values, as discussed in (Bruno et al. 2023). In statistical terms, RCA doesn’t satisfy the maximum likelihood condition, whereas entropy-based null models provide unbiased estimations (Garlaschelli & Loffredo 2008).
In this study, we employ the Inferred Comparative Advantage (ICA) to construct the $M_{c,i}$ matrix, comparing observed labour force with estimates from an entropy-based unbiased null-model which maintains fixed country-wide total labor and total sector strength. The null-model is known as the Bipartite Weighted Configuration Model (BiWCM) (Bruno et al. 2023). We adopt a similar methodology to RCA, adjusting for the accurate definition of the average country $E[L]_{c,i}$ by computing it using the BiWCM, such that the interpretation of the results is straightforward in the Economic Complexity framework. However, incorporating a null model through the BiWCM allows for more precise identification of comparative advantages with statistical significance. For instance, figure 1 illustrates the p-values of empirical values $L_{c,i}$ using a colour scale ranging from blue (indicating zero) to red (indicating one). Before feeding the ICA ($M_{c,i}$) matrix described in figure 1 into the EFC algorithm outlined in equation (1), we add an additional step to ensure comparability of Fitness and Complexity values across years.
The EFC algorithm computes a fixed point to determine the actual values of the Fitness and Complexity indices for each year independently. Consequently, in principle, it is not possible to compare the index values across different years due to the dynamic changes in the bipartite networks. To address this issue, we employ a technique, the "dummy trick", as elaborated in (Mazzilli et al. 2024). This involves augmenting the $M_{c,i}$ matrix with an additional row, referred to as the "dummy" row, which represents a country specialised in all industries. Thanks to the scale invariance property of the algorithm, we assign a constant Fitness value of 1 to this dummy country for every year. This approach establishes a fixed reference scale over time, enabling the comparison of Fitness and Complexity values across different years.
@@ 4.2 Econometric framework
It is now possible to compute the Industrial Complexity Index $Q_{i}$ used to construct our measure of Labour-Weighted Fitness, as described in equation 2, which in turn is decomposed into a within and a between term (equation 6). In particular, as mentioned in Section 2.2, we are interested in the decomposition term that captures the shift of labour shares across industries, that is, the between term. We use this term to explore its correlation with employment growth, wage inequality and the labour share of the economy, using the following specification:
@@
$g_{c,t}=\frac{emp_{c,t}}{emp_{c,t-1}}$ (9)
Wage inequality is constructed using the ILO Statistics on Labour Income and Inequality described in Section 3.3. In order to capture the effect of structural change on the distribution of income, we take three different measures of wage inequality: 9th to 1st percentile, 9th to 5th percentile, and 5th to 1st percentile:
@@
Finally, $X_{c,t}$ is a vector of time varying controls at the country level, including GDP per capita, population (from ARDECO), share of R&\&&D investments over GDP, share of exports over GDP (from the World Development Indicators).
All equations are estimated using OLS country ($\sigma_{c}$) and time ($\tau_{t}$) fixed effects, with errors clustered by country and year. Preliminary results of the descriptive econometric model are provided in the next section, along with an illustration of the results produced by the construction of the specialisation matrix and structural decompositions.
@@ 5.1 Descriptive results
We begin by commenting on the patterns exhibited by the specialisation matrix, which has been obtained by filtering the country-industry employment matrix for the year 2010 as described in Section 4.1. Figure 1 contains information on the statistical significance of individual country-industry linkages, with values going from more significant (red) to less significant (blue). The rows of the matrix represent countries, while the columns identify industries. Countries have been ordered from higher Fitness (top) to lower Fitness (bottom), and industries are sorted from the least (left) to the most (right) complex.
Figure 1: BIWCM Specialisation Matrix
At a first visual inspection, the matrix displays a pattern ascribable to the nested pattern (Bustos et al. 2012) that typically characterises country-product RCA matrices, filtered using the Balassa index (Patelli et al. 2023). Precisely, the distribution of the statistically significant linkages between countries and industries that describes countries’ diversification patterns has a triangular shape, which can be used as a broad approximation of the nested nature of the system being observed. While the concept of nestedness has originated in evolutionary biology, it has found vast application in the economic complexity literature: productive systems that exhibit a nested pattern reveal the presence of correlations within the system that cannot be accounted for using linear models or average approximations. In economic terms, nestedness indicates that industrial activities that are pursued by low-Fitness countries are also pursued by high-Fitness countries, which are however also more diversified.
@@
In the specific case described in figure 1, the economic interpretation of nestedness holds only in part; as a matter of fact, the top left corner – where high-Fitness countries and low-Complexity industries are located – appears to be less populated that one would expected in a fully nested matrix. While a similar pattern emerges also in other systems – such as the one of scientific production (Patelli et al. 2023) – in this case we find the observed result consistent with the process of structural change. When it comes to the employment dimension of the industrial production, high-Fitness countries tend to have higher diversification and to develop a comparative advantage in capital- or knowledge-intensive industries, at the expensive of more labour-intensive (and low-complexity) ones. As a result, industrialising countries with a comparative advantage in labour intensive activities (often due to lower wages) tend to specialise in the latter industries, and to have lower diversification. In substance, the sample of European countries seems to be divided in two subsamples of countries at different stages of structural transformation, with a minority of more mature countries and a majority of industrialising countries. The dynamic of Fitness rankings shows clearly which countries belong to each group, with the more mature economies being Germany, France, Italy, United Kingdom and Spain, followed by emerging countries such as Poland and Czechia.
Figure 2: Fitness rankings: 2010-2018
@@
We now move towards an analysis of the variation in LWF over the time period 2010–2018. Figure 3 shows the difference in LWF between 2010 and 2018, computed as per equation 6, with its respective between and within component. We shall recall that the former measures the contribution of structural change – defined here as the movement of labour to more complex industries – to the variation in LWF over the time period under consideration; while the latter picks up the change in LWF due to increasing complexity of the industries in which countries develop an (inferred) comparative advantage. Overall, figure 3 indicates that the between and within component are often anticorrelated. This can be read in economic terms by considering the reasonable assumption that as industries become more complex, they also become less labour intensive. The within component captures if the industries in which a country is specialised become more complex – for instance, as a result of technical change within the industry. By doing so, labour is shed by such complex industries, flowing into industries with lower complexity – a pattern associated with a between component of the opposite sign to the within component.
@@
To have a clearer idea of what happened in European countries over the 2010–2018, figure 4 describes the same quantities as figure 3, but for two different time periods: 2010–2014 (left panel) and 2014–2018 (right panel). The figure indicates that the first time period the variation in LWF fitness was negative for most countries, mostly driven by the decreasing within-industry complexity; it must be noted that this period coincides with the aftermaths of the 2008 financial crises, which has severely compromised economic activity in Europe across the board. In the following period, however, European countries show a much better performance, with most countries increasing their LWF, in many case due to within industry complexity and between industry reallocation of labour. One interesting case is the one of Poland, which is an outlier in both period: the country shows a huge within industry growth of complexity in 2010–2014, which is however followed by a negative between industry reallocation of labour in the following time period – hypothetically because of the lower labour intensity of its productive structure that has shifted towards increasingly knowledge-intensive activities.
Figure 3: Labour-weighted Fitness decomposition: 2010-2018
Figure 4: Labour-weighted Fitness decomposition: 2010-2014 and 2014-2018

### v2

@@ 2.1 The Economic Complexity framework
In order to measure the Fitness of countries and the Complexity of European industries, we rely on the Fitness and Complexity algorithm proposed by Tacchella et al. (2012):
[…]
@@
with $Q^{(0)}_{i}=1$. The algorithm takes as input a binary country-industry employment specialisation matrix $M_{c,i}$ constructed by cross-tabulating industrial employment data across European countries. The matrix is binarised according to the industrial specialisation profile of each country and indicating whether country $c$ displays a comparative advantage in industry ii. [^3] In this work, the binary specialisation matrix $M_{c,i}$ – which identifies industries in which countries are specialised – is obtained using an entropy maximising null model (the Bipartite Weighted Configuration Model by Bruno et al. (2023)) and implementing a method to ensure the comparability of Fitness and Complexity values over time. A more in-depth explanation of the construction of the matrix are explained in Section 4.1. In Appendix B.1 we detail the advantages of using an entropy-maximising null model to measure comparative advantages, over traditional methods such as the Revealed Comparative Advantage (Balassa, 1965).
Starting from this matrix, the iterative formula in Equation 1 jointly determines the values of country Fitness (first row of Equation 1) and industry complexity (second row of Equation 1). The intuition of the algorithm is straightforward: a country is assigned a high Fitness if it is specialised in many complex industries. Conversely, an industry is assigned a high Complexity if it hard to find, and when it is found, it appears in countries that are highly diversified. More formally, the Fitness $F_{c}$ of country $c$ at each time period corresponds to the weighted sum of the Complexities $Q_{i}$ of the industries i=1,,Ni=1,...,N in which the country holds a comparative advantage, as identified by the binary specialisation matrix $M_{c,i}$. In turn, the complexity $Q_{i}$ of industry $i$ corresponds to the inverse of its ubiquity weighted by the inverse of the Fitness of the countries which have that industry in their portfolio. More complex industries, therefore, are the least ubiquitous ones – proxying for the rareness of the capabilities required by those – considering also how ’fit’ are the countries that hold a specialisation in that industry. The algorithm is iterated $n$ times, until it converges. At each step, Fitness and Complexity values are normalised by dividing them by the respective average values. A feature that makes Fitness a suitable indicator to measure structural change as the process of capabilities accumulation structure is that it considers the cumulative nature of capabilities by considering explicitly the diversification of countries (weighted by the complexity of their industries). This feature is absent in other methods, such as the Economic Complexity Index (ECI) (Hausmann and Hidalgo, 2009), which do not consider diversification when measuring the complexity of countries’ productive structures (Cristelli et al., 2013; McNerney et al., 2025; Neffke et al., 2026). This property comes in handy to decompose the Economic Fitness measure and isolate its structural change component, as explained in the next subsection.
@@ 2.2 Labour-Weighted Fitness and decomposition
The starting point to construct a measure of structural change that captures the reallocation of labour across industries is to manipulate the Fitness equation (the first row of Equation 1) in such a way as to weight the sum of the complexities $Q_{i}$ by their labour share. Thus, for $N$ industries, the Labour-Weighted Fitness can be expressed as:
@@
Where $\theta_{i,t}$ is the employment share of industry $i$ at time $t$ for any given country. The resulting LWF is therefore an index of the Fitness of countries, where more weight is given to industries that are not only complex, but which also absorb a larger employment share, indicating the presence of knowledge-intensive capabilities across a larger portion of the workforce. To extract our measure of structural change, we now propose a structural decomposition of LWF inspired by the productivity decompositions adopted in the structuralist literature (McMillan et al., 2014, 2017; de Vries et al., 2015), which originates from the original work of Fabricant (1942). We decompose $\Delta LWF_{t}$ – the variation of LWF between time $t$ and tkt-k, with $0\leq k\leq T$ – in within- and between-sector components as follows:
@@
$(=)\ \underbrace{\sum_{i=1}^{N}\theta_{i,t-k}\Delta Q_{it}}_{\text{WITHIN TERM}}+\underbrace{\sum_{i=1}^{N}Q_{it}\Delta\theta_{i,t}}_{\text{BETWEEN TERM}}.$ (6)
Therefore, rearranging terms, Equation 3 yields the decomposition in the two distinct terms described in Equation 6:
  • $\sum_{i=1}^{N}\theta_{i,t-k}\Delta Q_{it}$ is the within-sector term of structural change (the contribution of increasing complexity at the sectoral level to the change in labour-weighted Fitness);
@@
The between-component, given by $\sum_{i=1}^{N}Q_{it}\Delta\theta_{i,t}$, identifies the measure of structural change of interest. This will be used to explain the observed trends in wage inequality, employment growth, and functional distribution of income.
@@ 3 Data
EUROSTAT’s Structural Business Survey.
The main data source used in this paper is the Eurostat Structural Business Statistics (SBS). The SBS describe the structure and performance of industrial sectors in the in European countries. They cover the ‘business economy’ (NACE Rev. 2 sections B to N and division 95) which includes: i) Industry; ii) Construction; iii) Distributive trades; and iv) Services. The data on industries is aggregated at the 4 digit level of the NACE classification, for a total of 218 NACE industries. SBS are based upon data for enterprises, which is obtained for national statistical offices and aggregated at the industry level based on the enterprise’s principal activity. Businesses in financial services (NACE Rev. 2 Section K) are excluded because of their specific characteristics and the unavailability of information in this industry, often due to confidentiality issues. In addition, SBS does not cover agriculture, forestry and fishing (NACE A), public administration (NACE O) and non-market services such as education (NACE P) and health (NACE Q).
As the analysis conducted here relies on patterns of industrial employment, we use the SBS data to describe the employment share of 4 digit NACE industries. To do so, we extract the SBS indicator that quantifies the number of persons employed [^4] Following the Eurostat’s SBS definition, this indicator ”is defined as the total number of persons who work in the observation unit (inclusive of working proprietors, partners working regularly in the unit and unpaid family workers), as well as persons who work outside the unit who belong to it and are paid by it (e.g. sales representatives, delivery personnel, repair and maintenance teams). It excludes manpower supplied to the unit by other enterprises, persons carrying out repair and maintenance work in the enquiry unit on behalf of other enterprises, as well as those on compulsory military service.” Source: https://ec.europa.eu/eurostat/cache/metadata/en/sbs_esms.htm#meta_update1663843213928. in each industry. In order to maximise data availability, we restrict the analysis to the 27 countries that were part of the European Union until 2018, observed between 2010 and 2018.
@@
ARDECO.
The Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy (ARDECO), [^5] The ARDECO dataset is available, free of charge, from the European Commission’s website: https://knowledge4policy.ec.europa.eu/territorial/ardeco-database_en. is a dataset curated and maintained by the Commission’s Joint Research Centre. It includes information on European countries and regions at different levels of disaggregation. While the dataset is based on the Eurostat’s SBS illustrated in the previous section, the data is complemented – where needed – with national and international sources. The data covers different topics, such as population, employment, labour costs, domestic product and capital formation. For the purpose of the analysis carried out in this paper, indicators on population, GDP per capita (constant prices), employment growth, compensation of employees (at constant prices) and gross value added (at constant prices) have been selected. The last two indicators are used to construct a measure of labour share of the economy, computed as the ratio of employee’s aggregate compensation over gross value added, in each country and year.
ILO Statistics on Labour Income and Inequality.
The data on wage inequality is obtained from the International Labour Organisation’s Statistics on Labour Income and Inequality. We focus on data regarding the labour income distribution, [^6] Available at: https://www.ilo.org/shinyapps/bulkexplorer46/?id=LAP_2LID_QTL_RT_A. which despite being based on modelled estimates, it provides information on average wages at all deciles of the labour income distribution across the countries in the sample, for the entire time period covered (2010 – 2018). In particular, we use information on labour income at the ninth, fifth and first deciles to produce wage ratios between ninth and first, fifth and first, and ninth and fifth deciles to capture the degree of inequality in labour income at different points of the distribution.
World Development Indicators.
@@ 4 Methods
4.1 From Inferred Comparative Advantages to a measure of Labour-weighted Fitness
A common starting point in the construction of Economic Complexity indicators is the observation of countries’ specialisation profiles across sectors or products. In practice, these profiles are typically identified through measures of comparative advantage, most notably Balassa’s (1965) Revealed Comparative Advantage – which defines specialisation of country $c$ in industry $i$ by comparing the observed relative weight of an activity in a given country to its expected value – that is, global average share of the typical country in the same activity. Specialisation is therefore defined by this share being larger than one. Nevertheless, this approach has a number of shortcomings: as pointed out by (Bruno et al., 2023), RCA is sensitive to extreme values and skewed distributions – like it is the case for industrial employment data – with outliers weighting disproportionately on the expected value used as denominator in the RCA formula. Moreover, RCA treats all country-activity pairs as independent, missing correlations that naturally arise in real networks.
In order to address these shortcomings, [^7] The issues noted above and described in Bruno et al. (2023) affect skewed data – like industrial employment data – to a larger extent as compared to activity data traditionally used in Economic Complexity, such as international trade data. in this paper we adopt a recent approach proposed by Bruno et al. (2023), the Inferred Comparative Advantage (ICA), to construct the binary country-sector matrix $M_{c,i}$. Rather than relying on ’raw’ RCA thresholds, ICA compares the observed allocation of labour across country-sector pairs with the distribution implied by an entropy-based unbiased null-model – the Bipartite Weighted Configuration Model (BiWCM) – that preserves two key aggregate features of the matrix, namely total employment at the country level and total employment at the sector level across all analysed countries. Further details on the methodology are provided in Appendix B.1.
Incorporating a null model allows us to identify specialisation patterns in an unbiased and more statistically robust way. For instance, Figure 1 reports the p-values associated to the empirical values of $M_{c,i}$ using a colour scale ranging from blue (zero), representing non significant country-sector links, to red (one), representing instead statistically significant links. Before feeding the ICA ($M_{c,i}$) matrix described in Figure 1 into the EFC algorithm outlined in Equation (1), we implement an additional step to ensure comparability of Fitness and Complexity values across years.
Since the EFC algorithm is computed separately for each year, the numerical values of Fitness depend on the structure of the underlying country–sector matrix and may therefore vary in scale over time. Raw Fitness values are thus not directly comparable across years. Following (Mazzilli et al., 2024), we address this issue by adding a dummy country specialised in all sectors and rescaling all Fitness values relative to this fixed benchmark, thereby ensuring temporal comparability. More details on this procedure are provided in Appendix B.2.
@@ 4.2 Econometric framework
It is now possible to compute the Industrial Complexity Index $Q_{i}$ used to construct our measure of Labour-Weighted Fitness, as described in Equation 2, which in turn is decomposed into a within and a between term (Equation 6). In particular, as mentioned in Section 2.2, we are interested in the decomposition term that captures the shift of labour shares across industries, that is, the between term, which captures the process of structural change isolated from the total change in labour-weighted Fitness. In other words, the between component of our decomposition pins down the changes in total fitness that derive from labour moving towards more complex industries, rather than industries becoming more complex themselves. We use this term to explore its correlation with employment growth, wage inequality and the labour share of the economy, using the following specification:
@@
$g_{c,t}=\frac{emp_{c,t}}{emp_{c,t-1}}.$ (8)
Wage inequality is constructed using the ILO Statistics on Labour Income and Inequality described in Section 3. In order to capture the effect of structural change on the distribution of income, we take three different measures of wage inequality: 9th to 1st percentile, 9th to 5th percentile, and 5th to 1st percentile:
@@
Finally, $X_{c,t}$ is a vector of time varying controls at the country level, including GDP per capita, population (from ARDECO), share of R&\&D investments over GDP, share of exports over GDP (from the World Development Indicators). All equations are estimated using OLS country ($\sigma_{c}$) and time ($\tau_{t}$) fixed effects, with errors clustered by country and year.
@@ 5.1 Descriptive results
We begin by commenting on the patterns revealed by the specialisation matrix, which has been obtained by filtering the country-industry employment matrix for the year 2010 as described in Section 4.1. Figure 1 contains information on the statistical significance of individual country-industry linkages, with values going from more significant (red) to less significant (blue). The rows of the matrix represent countries, while the columns identify industries. Countries have been ordered from higher Fitness (top) to lower Fitness (bottom), and industries are sorted from the least (left) to the most (right) complex.
Figure 1: ICA Specialisation Matrix
  • Notes: The figure reports the ICA specialisation matrix based on the BiWCM filtering procedure, evaluated for year 2018, where countries are displayed in the y-axis (higher fitness on the top) and the sectors in which they specialised in the x-axis (more complex industries to the right). Coloured cells indicate the statistical significance attributed to specialisation of a country in an industry. The colours on the colour bar report the (1 - pvalue) figures, with higher (statistically significant) values in red, and lower (non statistically significant) values in blue.
Rather, it points to a highly ordered structure in which diversification and specialization are systematically related. In economic terms, industries observed in low-Fitness countries are generally also observed in high-Fitness countries, whereas high-Fitness countries are distinguished by their ability to operate in many additional, less ubiquitous activities. At a first visual inspection, the matrix displays a nested structure (Bustos et al., 2012) – a typical feature of specialisation matrices (Patelli et al., 2023) indicating the presence of a hierarchical structure in the observed activities. Intuitively, nestedness means that the specialisation pattern of less diversified countries tends to form a subset of that of more diversified ones: industries in which less diversified economies are significantly present are typically also present in more diversified ones, while the latter are active in a wider range of industries. In the matrix, this appears as a roughly triangular pattern in the distribution of the statistically significant links between countries and industries. This type of structure, originally studied in evolutionary biology, has become central in the economic complexity literature (Neffke et al., 2026).
@@
In the case described in Figure 1, the economic interpretation of nestedness holds only in part; the top left corner – where high-Fitness countries and low-Complexity industries are located – appears to be less populated than expected in a fully nested matrix. While a similar pattern emerges also in other systems, such as in the specialisation matrix describing the scientific production of countries (Patelli et al., 2023), in this case we find the observed result consistent with the process of structural change. High-Fitness countries tend to have higher diversification and develop a comparative advantage in capital- or knowledge-intensive industries, at the expense of more labour-intensive (and low-complexity) ones. As a result, countries with comparative advantages in labour intensive activities (often reflecting lower labour costs) tend to specialise in those industries and exhibit lower levels of diversification. In broad terms, the sample of European countries we analyse appears to be split into two main groups at different stages of structural transformation: a smaller set of more mature, service-based economies and a larger group of less diversified countries. The evolution of Fitness rankings reported in Figure 2 clearly illustrates this divide. The upper part of the ranking is stably occupied by the more mature Western European economies – most notably Germany, France, Italy, United Kingdom and Spain – followed by a second group formed mainly by emerging Central and Eastern European (CEE) countries.
Figure 2: Fitness rankings (2010-2018).
  • Notes: The figure displays the evolution of Fitness country rankings based on ICA specialisation matrices between years 2010 and 2018. Countries are sorted in descending order by their Fitness in 2010 (left axis) and 2018 (right axis). Dots along the lines in the spaghetti plot identify years between 2010 and 2018.
@@
We now move towards an analysis of the variation in LWF over the time period 2010–2018. Figure 3 shows the difference in LWF between 2010 and 2018, computed as per Equation 6, with its respective between and within-component. We shall recall that the former measures the contribution of structural change – defined here as the movement of labour to more complex industries – to the variation in LWF over the time period under consideration; while the latter picks up the change in LWF due to increasing complexity of the industries in which countries develop a (inferred) comparative advantage. The barplot in Figure 3 indicates that the between- and within-component are often anticorrelated. This is consistent with the reasonable assumption that as industries become more complex, they also become less labour intensive. To make an example, the industries that exhibits the largest increase in complexity is Mining of Hard Coal – an industry that has become progressively more technologically intensive, and less able to absorb workers (Paredes and Fleming-Muñoz, 2021). The within-component captures if the industries in which a country is specialised become more complex – for instance, as a result of technological upgrading within the industry. By doing so, labour is shed by such complex industries, flowing into industries with lower complexity. This employment reallocation is thus often associated with a between-component of the opposite sign to the within-component.
@@
To have a clearer idea of what happened in European countries over the 2010–2018 period, Figure 4 describes the same quantities as Figure 3, but for two different time periods: 2010–2014 (left panel) and 2014–2018 (right panel). The figure indicates that in the first time period the variation in LWF fitness was negative for most countries, mostly driven by the decreasing within industry complexity (blue bars). It must be noted that this period coincides with the aftermaths of the 2008 financial crises, which has severely compromised economic activity in Europe across the board. In the following period, however, European countries show a much better performance, with most countries increasing their LWF, in many case due to within industry complexity and between industry reallocation of labour.
Poland represents a particularly interesting and illustrative case, which is an outlier in both periods. In 2010–2014, the country shows a very substantial within-industry complexity growth, which accounts for almost the entirety of the increase in labour-weighted Fitness. In the subsequent time period, instead, the within-component becomes negligible, while the between-component turns negative. This pattern points to a two-stage process: an initial phase of upgrading, in which complexity increased within existing sectors with little accompanying reallocation of labour across industries, followed by a phase in which these same activities displayed weaker labour absorption, leading employment to shift towards less complex sectors. In this sense, the negative between component may reflect the declining labour intensity of a productive structure increasingly oriented towards more complex and knowledge-intensive activities. This interpretation is consistent both with the earlier literature on FDI-led post-socialist restructuring, which emphasises the role of foreign-owned manufacturing plants in technological upgrading and structural transformation (Domański, 2003), and with more recent evidence documenting Poland’s capability accumulation, increasingly diversified productive structure, deeper integration into global value chains, and the growing importance of knowledge-intensive services in the national economy (OECD, 2025).
Figure 3: Labour-weighted Fitness decomposition (2010-2018).
  • Notes: For each country along the vertical axis, the plot reports the changes in Labour-weighted Fitness between 2010 and 2018 (red dot) and its between and within-components, respectively represented by the yellow and blue bars. Countries are sorted in descending order by magnitude of change in Labour-weighted Fitness.
Figure 4: Labour-weighted Fitness decomposition (2010-2014 and 2014-2018).
  • Notes: For each country along the vertical axis, the plot reports the changes in Labour-weighted Fitness (red dot) and its between- and within-component, respectively represented by the yellow and blue bars. The left panel refers to the period between 2010 and 2014, while the right panel refers to the period between 2014 and 2018. Countries are sorted in descending order by magnitude of change in Labour-weighted Fitness.